OpenStudy (anonymous):

Factor the polynomial completly. 4x^2 - x - 14 ?

OpenStudy (uri):

Now what is the number when adding or subracting 1 and multiplying 14?

OpenStudy (anonymous):

before that...u have to divide by 4 to clear coefficient of x^2 @uri

OpenStudy (anonymous):

so sum is -1/4 and product is -14/4=-7/2. hmmm

OpenStudy (anonymous):

i dont think there is none! @uri

OpenStudy (anonymous):

well... you can always use\[(-b \pm \sqrt{b ^{2} - 4ac})/2a\]

OpenStudy (uri):

was away sorry.

ganeshie8 (ganeshie8):

you may try the uri way of factoring which is simple 14x4 = 56 so you need to find two numbers that multiply to 56, and add up to -1

OpenStudy (anonymous):

-8 and 7 ??

ganeshie8 (ganeshie8):

can you find them ? if not you can use the quadratic formula directly

OpenStudy (anonymous):

-8 and 7 @ganeshie8 and @uri

ganeshie8 (ganeshie8):

yes ! thats right !! we rewrite the given polynomial

OpenStudy (anonymous):

i think thats wrong mate...doen't fit the equaton

OpenStudy (uri):

4x^2 - x - 14 4x^2-8x+7x-14 4x(x-2)+7(x-2)

OpenStudy (anonymous):

but isn't the equation supposed to reduce tto zero when we put value of roots??

ganeshie8 (ganeshie8):

it wil reduce to 0, this is simply oly :)

ganeshie8 (ganeshie8):

*simple

OpenStudy (uri):

So answer is (4x+7)(x-2)

ganeshie8 (ganeshie8):

looks right to me

OpenStudy (anonymous):

can you tell me how it will be zero?? @ganeshie8

OpenStudy (uri):

0?

ganeshie8 (ganeshie8):

wat are the roots we got here ha ? x-2 = 0 x=2 plugin x =2 in given polynomial wat do we get ?

OpenStudy (anonymous):

@uri

OpenStudy (uri):

We have to factorize this eq by finding the GCF..

ganeshie8 (ganeshie8):

repeat the same thing for other factor 4x+7 = 0 x = -7/4 plugin this value in given polynomial and see wat u get

OpenStudy (anonymous):

get it now..thx guys.:) btw wats GCF

OpenStudy (uri):

GREATEST COMMON FACTOR.

ganeshie8 (ganeshie8):

great :) GCF = Greatest Common Gactor we have used it when grouping things

OpenStudy (uri):

we expand the eq.

OpenStudy (anonymous):

so whats my answer @uri and @ganeshie8

OpenStudy (uri):

(4x+7)(x-2)